OpenStudy (anonymous):

2/3 (x-4) = 2x How do you solve that ?

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{3}(x-4)=2x?\)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

\[\frac{2}{3}(x-4)=2x?\] right?

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

i hate fractions

OpenStudy (misty1212):

multiply both sides by 3 and get \[2(x-4)=6x\] is a good start

OpenStudy (anonymous):

I get everything except the 2x part

OpenStudy (anonymous):

Multiply the 2/3 by x/1

OpenStudy (anonymous):

What do you mean ?

OpenStudy (misty1212):

your fraction in front is \(\frac{2}{3}\) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Make the -4 12/3 and keep that 2/3x

OpenStudy (anonymous):

And make the 2x , 6/3x, it'll make it all Easter if it's all fractions

OpenStudy (misty1212):

lets go slow

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{3}(x-4)=2x\implies \cfrac{2(x-4)}{{\color{brown}{ 3}}}=2x \\ \quad \\ 2(x-4)={\color{brown}{ 3}}(2x)\implies 2x-8=6x\implies ?=x\)

OpenStudy (anonymous):

want a picture of wavy I'd do ?

OpenStudy (anonymous):

I don't wanna push you guys but can you do the problem step-by-step because im not getting anything? It would be helpful.

OpenStudy (anonymous):

Sure retro

OpenStudy (anonymous):

Thanks jdoe

OpenStudy (anonymous):

I wanna see everyone of your answer with your ways and see which one easiest so I can do them on my own later :D

OpenStudy (misty1212):

\[\frac{2}{3}(x-4)=2x\] the annoying part is the fraction \(\frac{2}{3}\) we can get rid of the denominator by multiplying both sides by 3

OpenStudy (anonymous):

Okay so 2x would be 6x....

OpenStudy (misty1212):

that will get rid of the fraction \[3\times \frac{2}{3}(x-4)=3\times 2x\\ 2(x-4)=6x\]

OpenStudy (anonymous):

Won't let me post picture

OpenStudy (misty1212):

then it should be more or less routine distribute first get \[2x-8=6x\]

OpenStudy (misty1212):

subtract \(2x\) from both sides to get \[-8=4x\]

OpenStudy (anonymous):

I got the answer

OpenStudy (misty1212):

then divide both sides by \(4\)

OpenStudy (anonymous):

X= -3

OpenStudy (anonymous):

You did it wrong ......

OpenStudy (misty1212):

actually no it is not \(-3\)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I wished it would let me show you my work

OpenStudy (anonymous):

x=-2 @misty1212 ?

OpenStudy (anonymous):

Why won't it let me put pics ?

OpenStudy (anonymous):

@misty1212